OpenStudy (shelby1290):

To solve the following linear system by elimination, Brent first multiplied each equation by 10. Explain why he did this step. Complete the solution. 0.3x - 0.5y = 1.2 0.7x - 0.2y = -0.1

OpenStudy (misty1212):

HI!!

OpenStudy (shelby1290):

Hey

OpenStudy (misty1212):

multiply by 10 get rid of annoying decimals

OpenStudy (misty1212):

\[0.3x - 0.5y = 1.2\\ 0.7x - 0.2y = -0.1\] \[3x - 5y = 12\\ 7x - 2y = -1\]

OpenStudy (shelby1290):

so are you saying that he multiplied 10 by 0.3....multiplied 10 by 0.5 and multiplied 10 by 1.2?

OpenStudy (misty1212):

yeah if u multiply one part by 10, u gotta multiply all by 10

OpenStudy (shelby1290):

Okay so the answer is: Brent multiplied by 10 to get rid of the decimal numbers

OpenStudy (misty1212):

yes

OpenStudy (doc.brown):

\[1+1+1=3\]\[10+10+10=3\times 10\]

OpenStudy (doc.brown):

The numbers are still there, I think Brent multiplied by 10 to get rid of the decimal places.

OpenStudy (shelby1290):

Okay should i add in what misty said before about how ... you'd get 3x -5y = 12

OpenStudy (shelby1290):

@doc.brown ^

OpenStudy (misty1212):

no

OpenStudy (misty1212):

\[3x - 5y = 12\\ 7x - 2y = -1\] u still have 2 do more work

OpenStudy (shelby1290):

okay so is there anything that i should calculate at all?

OpenStudy (misty1212):

multiply the first by 2, the second by -5 and add the y terms will go bye bye

OpenStudy (doc.brown):

Only if you want to follow the instructions, which say complete the solution.

OpenStudy (shelby1290):

@misty1212 and after doing that I'll find the solution, right?

OpenStudy (shelby1290):

If i multiply the first equation by 2 and the second by -5 and add the y terms

OpenStudy (shelby1290):

I'm still a little confused on what I'm supposed to write down

OpenStudy (doc.brown):

You want to know what x is. One way, is to divide the first one by 3.

OpenStudy (shelby1290):

0.3 divided by 3= 0.1 0.5 divided by 3= 0.16 1.2 divided by 3=0.4

OpenStudy (doc.brown):

Ah, sorry, I thought you were starting from\[3x - 5y = 12\\ 7x - 2y = -1\]

OpenStudy (shelby1290):

oops

OpenStudy (shelby1290):

3 divided by 3= 1 5 divided by 3=1.6 12 divided by 3=4

OpenStudy (doc.brown):

I'd do this, it's less work.\[3x - 5y = 12\\\frac{3x}{3} - \frac{5y}{3} = \frac{12}{3}\\\frac{\cancel{3}x}{\cancel{3}} - \frac{5y}{3} = \frac{12}{3}\\x-\frac{5}{3}y=4\\x=4+\frac{5}{3}y\]

OpenStudy (shelby1290):

thank you what's x and y

OpenStudy (doc.brown):

I don't know, stick x into the second equation.

OpenStudy (shelby1290):

did you find it already?

OpenStudy (shelby1290):

or is that what i'm supposed to do but btw i'm not exactly sure how to solve everything from here. Fractions confuse me.

OpenStudy (doc.brown):

No they don't, your teacher confused you. @misty1212 was having you make the y's equal and subtract them to get x by itself. I was getting x alone on one side and substituting it in to the other equation to get y by itself. Do you have a preference?

OpenStudy (shelby1290):

Nah but I'll just follow what you're doing and get x by itself

OpenStudy (shelby1290):

I'll do whatever it takes to get the correct answer!

OpenStudy (shelby1290):

so was misty's answer wrong?

OpenStudy (doc.brown):

They all work, from the first equation I got\[\color{red}{x=4+\frac{5}{3}y}\]Stick that in here\[7\color{red}{x}-2y=-1\]

OpenStudy (shelby1290):

7(4 + 5/3y) -2y = -1

OpenStudy (doc.brown):

Distribute the 7. 7(4)+7(5y/3)...

OpenStudy (shelby1290):

28+ 35y/3

OpenStudy (doc.brown):

\[28+\frac{35}{3}y-2y=-1\]Now what?

OpenStudy (shelby1290):

do i bring 28 to the other side?

OpenStudy (shelby1290):

then it would become -1-28

OpenStudy (doc.brown):

yup, then add the ys

OpenStudy (shelby1290):

okay so 37/3 ?

OpenStudy (shelby1290):

hold on, shouldn't i be subtracting not adding?

OpenStudy (shelby1290):

then the answer is: 33/3y=-29

OpenStudy (shelby1290):

@doc.brown ^

OpenStudy (doc.brown):

Can you imagine a half?

OpenStudy (shelby1290):

why, what do you mean?

OpenStudy (doc.brown):

What happens if you add a quarter to a half, what do you get?

OpenStudy (shelby1290):

3/4

OpenStudy (doc.brown):

That's right, the bottom number has to be the same and you add up the tops.

OpenStudy (doc.brown):

\[\frac{35}{3}y-\frac{6}{3}y\]

OpenStudy (shelby1290):

@doc.brown 29/3

OpenStudy (doc.brown):

\[\frac{29}{3}y=-29\]

OpenStudy (shelby1290):

yes

OpenStudy (shelby1290):

@doc.brown what's the next step

OpenStudy (shelby1290):

divide both sides but by what

OpenStudy (doc.brown):

You want the y to be alone, multiply each side by 3/29

OpenStudy (shelby1290):

1y=-3

OpenStudy (doc.brown):

Good, so if we know \(y=-3\) What's\[x=4+\frac{5}{3}y\]

OpenStudy (shelby1290):

x=4+5/3(-3) =-1 so the answer is -1 @doc.brown

OpenStudy (doc.brown):

Does it work, does \(x=-1\) and \(y=-3\) if\[0.3x - 0.5y = 1.2\]

OpenStudy (shelby1290):

0.3(-1) - 0.5 (-3) =1.2 Yea it works

OpenStudy (shelby1290):

@doc.brown Thank you so much for all the help!